# ExGRPO: Learning to Reason from Experience

### v1

@@ E.4 Ablation and Additional Results
E.5 Data Utilization
@@ Discussion.
u Code v Code u Code v Code $\Delta$
@@
Case #3 of Snowball Effects Question: ## [ Auxiliary area. The area helps to solve the problem ] Pythagorean Theorem (direct and inverse) In triangle ABCABC, heights AEAE and $CD$ are drawn. Find the side ABAB, if $BD=18,BC=30,AE=20$. Ground Truth: $\boxed{25}$ Response: (Low Entropy) Reward: 1.0Entropy: 0.0198CoT Validity (Judged by Qwen3-32B): Correct. To solve the problem, we need to use the properties of right triangles and the Pythagorean theorem. Here’s the step-by-step reasoning: 1. Identify the given information: $BD=18$ $BC=30$ $AE=20$ CDCD is a height from $C$ to ABAB, so $CD\perp AB$ AEAE is a height from AA to BCBC, so $AE\perp BC$ 2. Determine the position of point $D$: Since $D$ is the foot of the perpendicular from $C$ to ABAB, $BD=18$ and $BC=30$, we can find $AD$ using the Pythagorean theorem in triangle BCDBCD: $CD=\sqrt{BC^{2}-BD^{2}}=\sqrt{30^{2}-18^{2}}=\sqrt{900-324}=\sqrt{576}=24$ So, $CD=24$. 3. Use the area of the triangle: The area of triangle ABCABC can be calculated in two ways: Using $AE$ and $BC$: $\text{Area}=\frac{1}{2}\times BC\times AE=\frac{1}{2}\times 30\times 20=300$ Using $CD$ and $AB$: $\text{Area}=\frac{1}{2}\times AB\times CD=\frac{1}{2}\times AB\times 24$ Equating the two expressions for the area, we get: $\frac{1}{2}\times AB\times 24=300\implies AB\times 24=600\implies AB=\frac{600}{24}=25$ So, the length of $AB$ is $25$. </think> The final answer is $\boxed{25}$. Response: (High Entropy) Reward: 1.0Entropy: 0.3002CoT Validity (Judged by Qwen3-32B): Wrong. To solve the problem, we need to use the properties of the triangle and the Pythagorean theorem. Here is the step-by-step thought process: 1. Understand the Problem: We need to find the side ABAB of triangle ABCABC. We know the length of $BD=18$ and $BC=30$. We know the height $AE=20$. 2. Apply the Right Triangle Properties: Since $AE$ is a height, AEAE is perpendicular to BCBC, so triangle $AEB$ and triangle AECAEC are right triangles. We can use the Pythagorean theorem in triangles ABDABD and $AEC$. 3. Using the Pythagorean Theorem in $\triangle ABD$: In $\triangle ABD$, ADAD can be found using the Pythagorean theorem: $AD=\sqrt{AB^{2}-BD^{2}}$. 4. Using the Pythagorean Theorem in $\triangle AEC$: In $\triangle AEC$, ECEC can be found using the Pythagorean theorem: $EC=\sqrt{AC^{2}-AE^{2}}$. Since $BC=BD+DC$ and $D$ is on BCBC, DC=BCBD=3018=12DC=BC-BD=30-18=12. In $\triangle AEC$, $AC=EC$, because AEAE is the height from AA to $BC$. 5. Using the Area Equivalence: The area of the triangle can be calculated in two ways: Using base $BC$ and height AEAE: $\text{Area}=\frac{1}{2}\times BC\times AE=\frac{1}{2}\times 30\times 20=300$. Using base $AB$ and height CDCD: $\text{Area}=\frac{1}{2}\times AB\times CD$. 6. Find CDCD: From the area equivalence, $\frac{1}{2}\times AB\times CD=300$, so $CD=\frac{600}{AB}$. 7. Using the Pythagorean Theorem in $\triangle ACD$: In $\triangle ACD$, $AC=\sqrt{AD^{2}+CD^{2}}=\sqrt{(AB^{2}-BD^{2})+CD^{2}}$. Let’s use these steps to set up the equations and solve for ABAB. </think> Using the relationships and the Pythagorean theorem, we can derive the value of $AB$. Let’s first find $AB$ using the properties of the triangle and the given values. We know: $BD=18,BC=30,AE=20$ First, we calculate $CD$: $CD=\frac{600}{AB}$ Now, we use the Pythagorean theorem in $\triangle ABD$: $AD=\sqrt{AB^{2}-18^{2}}$ And in $\triangle AEC$: $AC=\sqrt{12^{2}+20^{2}}=\sqrt{144+400}=\sqrt{544}=4\sqrt{34}$ The area using base $BC$ and height $AE$ is: $\text{Area}=\frac{1}{2}\times 30\times 20=300$ The area using base $AB$ and height $CD$ is: $\text{Area}=\frac{1}{2}\times AB\times CD=\frac{1}{2}\times AB\times\frac{600}{AB}=300$ Thus, $CD=\frac{600}{AB}$ We now use the Pythagorean theorem in $\triangle ABD$ and $\triangle AEC$: $AD=\sqrt{AB^{2}-18^{2}}=\sqrt{AB^{2}-324}$ $AC=\sqrt{12^{2}+20^{2}}=4\sqrt{34}$ Since $AC$ is the hypotenuse of $\triangle AEC$: $AC=AE^{2}+EC^{2}$ We can now solve for $AB$ by equating the areas and solving the equations. Let’s now calculate the value of $AB$ using Python to find the accurate result.

### v2

@@ E.4 Ablation and Additional Results
Effects of Reintroducing Retired Queries.
We investigate whether periodically replaying previously solved (100%-success) queries can mitigate potential skill forgetting. Specifically, we reintroduce 10% of retired examples into the training batch every 25 steps. Contrary to intuition, this replay mechanism leads to a drop in both in-distribution and out-of-distribution performance, as shown in Table 7, suggesting that revisiting fully solved queries injects limited learning signal while displacing more informative mid-difficulty cases.
Table 7: Ablation on reintroducing retired queries. Bold indicates the better results.
In-Distribution Performance Out-of-Distribution Performance
AIME24 AIME25 AMC MATH Minerva Olympiad Avg. ARC-c GPQA MMLU-P Avg.
ExGRPO 31.6 18.7 66.3 87.4 36.0 50.1 48.4 84.7 37.4 52.9 58.3
[…]
E.5 Replay Memory and Computational Overhead
Table 8: Replay memory and time-per-step analysis. Times in seconds.
Model Method RAM Peak Step (s) Min. Step (s) Avg. Step (s)
Qwen2.5-Math-7B On-Policy 211.80 139.83 155.38
ExGRPO 1.77GB 256.00 151.36 184.92
Overhead +1.77GB +20.9% +8.2% +19.0%
Qwen2.5-Math-1.5B On-Policy 254.58 150.55 185.00
ExGRPO 1.05GB 228.90 163.92 188.85
Overhead +1.05GB -10.1% +8.9% +2.1%
The replay buffer introduces additional memory usage. In our largest setup (Qwen2.5-Math-7B on 8 GPUs), a buffer containing 202,011 trajectories requires approximately 1.77 GB of RAM at the end of training. This footprint is modest for modern hardware. Regarding computational overhead, ExGRPO is efficient because we do not recompute entropy for the entire buffer at each step. Instead, for each query sampled in a minibatch, we compute entropy only for its associated successful trajectories, typically a few dozen candidates. Empirically, as shown in Table 8, ExGRPO introduces acceptable overhead while providing substantial gains in reasoning performance.
E.6 Data Utilization
@@ Discussion.
Code Code Code Code $\Delta$
@@
Case #3 of Snowball Effects Question: ## [ Auxiliary area. The area helps to solve the problem ] Pythagorean Theorem (direct and inverse) In triangle ABCABC, heights AEAE and $CD$ are drawn. Find the side ABAB, if $BD=18,BC=30,AE=20$. Ground Truth: $\boxed{25}$ Response: (Low Entropy) Reward: 1.0Entropy: 0.0198CoT Validity (Judged by Qwen3-32B): Correct. To solve the problem, we need to use the properties of right triangles and the Pythagorean theorem. Here’s the step-by-step reasoning: 1. Identify the given information: $BD=18$ $BC=30$ $AE=20$ CDCD is a height from $C$ to $AB$, so $CD\perp AB$ AEAE is a height from AA to $BC$, so $AE\perp BC$ 2. Determine the position of point $D$: Since $D$ is the foot of the perpendicular from $C$ to $AB$, $BD=18$ and $BC=30$, we can find $AD$ using the Pythagorean theorem in triangle BCDBCD: $CD=\sqrt{BC^{2}-BD^{2}}=\sqrt{30^{2}-18^{2}}=\sqrt{900-324}=\sqrt{576}=24$ So, $CD=24$. 3. Use the area of the triangle: The area of triangle ABCABC can be calculated in two ways: Using $AE$ and $BC$: $\text{Area}=\frac{1}{2}\times BC\times AE=\frac{1}{2}\times 30\times 20=300$ Using $CD$ and $AB$: $\text{Area}=\frac{1}{2}\times AB\times CD=\frac{1}{2}\times AB\times 24$ Equating the two expressions for the area, we get: $\frac{1}{2}\times AB\times 24=300\implies AB\times 24=600\implies AB=\frac{600}{24}=25$ So, the length of $AB$ is $25$. </think> The final answer is $\boxed{25}$. Response: (High Entropy) Reward: 1.0Entropy: 0.3002CoT Validity (Judged by Qwen3-32B): Wrong. To solve the problem, we need to use the properties of the triangle and the Pythagorean theorem. Here is the step-by-step thought process: 1. Understand the Problem: We need to find the side ABAB of triangle ABCABC. We know the length of $BD=18$ and $BC=30$. We know the height $AE=20$. 2. Apply the Right Triangle Properties: Since $AE$ is a height, AEAE is perpendicular to $BC$, so triangle $AEB$ and triangle AECAEC are right triangles. We can use the Pythagorean theorem in triangles ABDABD and $AEC$. 3. Using the Pythagorean Theorem in $\triangle ABD$: In $\triangle ABD$, ADAD can be found using the Pythagorean theorem: $AD=\sqrt{AB^{2}-BD^{2}}$. 4. Using the Pythagorean Theorem in $\triangle AEC$: In $\triangle AEC$, ECEC can be found using the Pythagorean theorem: $EC=\sqrt{AC^{2}-AE^{2}}$. Since $BC=BD+DC$ and $D$ is on BCBC, DC=BCBD=3018=12DC=BC-BD=30-18=12. In $\triangle AEC$, $AC=EC$, because AEAE is the height from AA to $BC$. 5. Using the Area Equivalence: The area of the triangle can be calculated in two ways: Using base $BC$ and height AEAE: $\text{Area}=\frac{1}{2}\times BC\times AE=\frac{1}{2}\times 30\times 20=300$. Using base $AB$ and height CDCD: $\text{Area}=\frac{1}{2}\times AB\times CD$. 6. Find CDCD: From the area equivalence, $\frac{1}{2}\times AB\times CD=300$, so $CD=\frac{600}{AB}$. 7. Using the Pythagorean Theorem in $\triangle ACD$: In $\triangle ACD$, $AC=\sqrt{AD^{2}+CD^{2}}=\sqrt{(AB^{2}-BD^{2})+CD^{2}}$. Let’s use these steps to set up the equations and solve for ABAB. </think> Using the relationships and the Pythagorean theorem, we can derive the value of $AB$. Let’s first find $AB$ using the properties of the triangle and the given values. We know: $BD=18,BC=30,AE=20$ First, we calculate $CD$: $CD=\frac{600}{AB}$ Now, we use the Pythagorean theorem in $\triangle ABD$: $AD=\sqrt{AB^{2}-18^{2}}$ And in $\triangle AEC$: $AC=\sqrt{12^{2}+20^{2}}=\sqrt{144+400}=\sqrt{544}=4\sqrt{34}$ The area using base $BC$ and height $AE$ is: $\text{Area}=\frac{1}{2}\times 30\times 20=300$ The area using base $AB$ and height $CD$ is: $\text{Area}=\frac{1}{2}\times AB\times CD=\frac{1}{2}\times AB\times\frac{600}{AB}=300$ Thus, $CD=\frac{600}{AB}$ We now use the Pythagorean theorem in $\triangle ABD$ and $\triangle AEC$: $AD=\sqrt{AB^{2}-18^{2}}=\sqrt{AB^{2}-324}$ $AC=\sqrt{12^{2}+20^{2}}=4\sqrt{34}$ Since $AC$ is the hypotenuse of $\triangle AEC$: $AC=AE^{2}+EC^{2}$ We can now solve for $AB$ by equating the areas and solving the equations. Let’s now calculate the value of $AB$ using Python to find the accurate result.